# Active Learning of A Crystal Plasticity Flow Rule From Discrete Dislocation Dynamics Simulations

### v1

@@ 2.3 Additional constraints
The candidate stress component magnitudes are bound from above by Pa and their resulting global stress tensor is subject to a minimum eigen value (or eigen stress) constraint which ensures that the maximum magnitude of the eigenvalues of any applied stress tensor is at least Pa. This constraint was chosen to ensure that the load applied is sufficient to produce usable measurements of plastic distortion and strain rate within a reasonable amount of compute time. A more concrete exploration of distributions over the domain imparted by the constraints of the acquisition function is presented in Appendix A.
@@ 2.4 Discrete Dislocation Dynamics
Discrete dislocation dynamics simulations were performed using MoDELib [25] with a parameterization for -Fe at 320 K similar to ref. [26]. Simulation parameters influential to the characteristics of our results include the use of periodic boundaries and a system box size of approximately 1.5 m per side, adaptive time steps with subcycling, and omission of cross-slip. The mobility law employed which determines dislocation segment velocity incorporates non-Schmid effects and imparts a sensitivity to the component of the applied stress tensor when specified in a slip system aligned basis [27]. As illustrated in Fig. 9, the crystal has been rotated within the system box about the direction by 30 degrees to reduce the likelihood of annihilation of dislocations due to the periodic boundaries.
@@
After instantiating a targeted microstructure, batches of discrete dislocation dynamics simulations were performed under constant applied stress. To provide an interface between the DDD software written in C++ and the Python Bayesian optimization routines, a Python interface to MoDELib was implemented using pybind11 [28]. The python interface provides functionality to specify and instantiate various microstructures, as well as to run DDD while measuring slip system resolved quantities of interest. Specific to our purpose was the need to measure the plastic strain per slip system , stress tensor , and all slip system dislocation densities .
@@ 2.5 Processing of measurements
Second, reactions among dislocations and details of numerical implementation introduce statistical noise on the measured quantities. To reduce the influence of these noise sources, we calculate the strain rate as the slope of a line fit to clusters of 100 consecutive measurements of plastic distortion, as counted from the end of the simulation. Corresponding values of slip system densities , total density , and time are obtained by averaging over the 100 consecutive measurements.
@@ 2.6 Workflow and automation
The overall flow of data described by the preceding subsections is illustrated in Fig. 1. The 1-D diagram labeled as the GP model is a simple representation of Gaussian process regression and should not be mistaken for the actual model which is constructed over an 8-D domain. Management and automation of the outermost active learning loop was implemented using Python, within which Parsl [29] was used to automate interactions with the SLURM workload manager in a way which minimized human interaction. To ensure the sanity and productivity of each batch of simulations, stopping points were inserted immediately prior to instantiating a batch of simulations, as well as after each batch of simulations concluded. Because a SLURM allocation was only allotted a restricted number of hours, Python and Parsl were also used to automatically restart each simulation after an allocation expired for a total of 5 allocations of 16 hours per simulation in a batch of 10 simulations per iteration of the outermost loop.
@@ 3.1 Discrete dislocation dynamics simulations
Evolution of a typical DDD simulation is illustrated in Fig. 2 at three stages: (a) instantiation of 1 m radius shear loops and prismatic loops as described in section 2.4, (b) evolution until reaching 0.2% strain, and (c) final higher density configuration. These images correspond to the simulation displayed in Fig. 3a.
@@
The prominence of rising densities is a consequence of our use of periodic boundaries. In a DDD convention in which plastic distortion of a shear loop resolved on its own slip system is always negative, applying a positive resolved shear stress causes loops within that slip system to shrink, contributing positively to the strain rate, while a negative will cause them to grow and contribute negatively to the slip system’s strain rate. Thus the dislocation loops which contribute positively to strain rates are short-lived, shrinking until their inevitable self annihilation and often do not persist far beyond the 0.2% strain threshold. For this reason the data sets produced by our simulations are sparse in the regions of positive strain rates. Fig. 3c demonstrates an occasional example counter to this trend, in which the applied load and configuration led to a slow shrinking of the majority of dislocation loops.
@@
Fig. 3 illustrates several other notable features found within our measurements. Deformation within the simulation presented in the second column is dominated by a single slip system whose dislocation density and plastic deformation show consistent, if not accelerating, growth. This is in contrast to the simulation of the first column in which the dislocation densities of multiple slip systems are growing at rates which vary in time, due to interactions between dislocations on different slip systems. It could be argued that the initially high strain rate of the first slip system was due to lack of interaction with the other slip systems resulting in a period of free movement which was soon to be suppressed when interactions became prominent, producing varying strain rates and noise in the functional relationship . The leading theory of how these slip system interactions reduce strain rates is via the formation of junctions between dislocations whose mobilities are reduced or negligible [30, 31]. It should be noted that the specific configurational information which induces these interactions and the resulting strain rates is lost when the configuration is represented as dislocation densities per slip system, but could be recovered from the raw logs of the DDD simulations. The significant variations in strain rate measurements for a given stress tensor and list of densities are then represented by the fitted noise hyperparameter of the GP model.
@@ 3.2 Model evolution
The non-negligible magnitude of the noise in DDD strain-rate measurements induces a lower limit on the error of our model, forcing the average error of its predictions to be at least as bad as the noise. This is shown by the similarity between the curves labeled ‘AVG of STDEV of test data’ and ‘AVG error of predictive mean’ in Figure 4a. This echos observations by another study[6] which discussed the effect of noise on fitted coefficients of the Kocks-Mecking model using a large database of DDD simulations. We concur that methods of reducing and compensating for the noise obtained in these measurements are needed to improve model predictions. While optimizing the reduced dimensional representation of all 24 slip system densities could possibly reduce noise, e.g. forest[32], coplanar[7, 33], or glissile/sessile, our choice of a simpler specification avoided further complicating the acquisition function.
@@ Appendix A Distributions of simulation targets and eventual measurements
In Section 2.2 we described how our acquisition function was designed to address the differences in dimensionality between model domain and simulation parameters, while in Sections 2.3 and 2.5, we described simulation constraints and transformations, respectively, intended to improve the distributions of data presented to the model in the spirit of regularization so that it may produce a better fit. To assess the influence of these designs, transformations, and constraints we present here a study of the intrinsic distributions they collectively impart. We compare these intrinsic distributions to the final collection of training data obtained from measurements to gain insight into the ability of our BO implementation to navigate parameter space under the influence of the intrinsic distributions.
@@
The distribution of BO targets for particular slip system densities resemble neither distribution of the first or the remaining slip system particularly well, indicating that the acquisition function had greater liberty in guiding slip system density specifications. However, the distribution of measured slip system density values which evolved during the simulations more closely resemble the distribution labeled “random, all but first slip system,” with the exception of the presence of a large number of slip systems having dislocation density. So while the acquisition function had the most liberty to determine dislocation density targets, their evolution during the simulation caused measurements to drift significantly far from their targets.

### v2

@@ 2.5 Additional constraints
The candidate stress component magnitudes are bound from above by Pa and their resulting global stress tensor is subject to a minimum eigen value (or eigen stress) constraint which ensures that the maximum magnitude of the eigenvalues of any applied stress tensor is at least Pa. This constraint was chosen to ensure that the load applied is sufficient to produce usable measurements of plastic distortion and strain rate within a reasonable amount of compute time. While the constraints on applied stresses and dislocation densities limit the acquired data to relatively high driving forces they do not result in any loss of generality in the computational workflow and can be easily relaxed with more computational resources to target regimes with low driving forces. A more concrete exploration of distributions over the domain imparted by the constraints of the acquisition function is presented in Appendix A.
@@ 2.6 Discrete Dislocation Dynamics
Discrete dislocation dynamics simulations were performed using MoDELib [po2014recent] with a parameterization for -Fe at 320 K similar to ref. [MAHLER2021100814]. Simulation parameters influential to the characteristics of our results include the use of periodic boundaries and a system box size of approximately 1.5 m per side, adaptive time steps with subcycling, and omission of cross-slip. The mobility law employed which determines dislocation segment velocity incorporates non-Schmid effects and imparts a sensitivity to the component of the applied stress tensor when specified in a slip system aligned basis [PO2016123]. As illustrated in Fig. 9, the crystal has been rotated within the system box about the direction by 30 degrees to reduce the likelihood of annihilation of dislocations due to the periodic boundaries.
@@
After instantiating a targeted microstructure, batches of discrete dislocation dynamics simulations were performed under constant applied stress. To provide an interface between the DDD software written in C++ and the Python Bayesian optimization routines, a Python interface to MoDELib was implemented using pybind11 [pybind11]. The python interface provides functionality to specify and instantiate various microstructures, as well as to run DDD while measuring slip system resolved quantities of interest. Specific to our purpose was the need to measure the plastic strain per slip system , stress tensor , and all slip system dislocation densities .
@@ 2.7 Processing of measurements
Second, reactions among dislocations and details of numerical implementation introduce statistical noise on the measured quantities. To reduce the influence of these noise sources, we calculate the strain rate as the slope of a line fit to clusters of 100 consecutive measurements of plastic distortion, as counted from the end of the simulation. Corresponding values of slip system densities , total density , and time are obtained by averaging over the 100 consecutive measurements. An additional consequence of these noise sources was the infrequent presence of outlying data in which changes to plastic distortion were in the opposite direction of the applied load. In the context of producing a flow rule for CP, these data points are considered unproductive deviations from the physics which it aims to capture. To reduce the influence of this noise on the quality of the model, the strain rates of any measurements having were omitted from the training data set.
@@ 2.8 Workflow and automation
The overall flow of data described by the preceding subsections is illustrated in Fig. 1. The 1-D diagram labeled as the GP model is a simple representation of Gaussian process regression and should not be mistaken for the actual model which is constructed over an 8-D domain. Management and automation of the outermost active learning loop was implemented using Python, within which Parsl [10.1145/3307681.3325400] was used to automate interactions with the SLURM workload manager in a way which minimized human interaction. To ensure the sanity and productivity of each batch of simulations, stopping points were inserted immediately prior to instantiating a batch of simulations, as well as after each batch of simulations concluded. Because a SLURM allocation was only allotted a restricted number of hours, Python and Parsl were also used to automatically restart each simulation after an allocation expired for a total of 5 allocations of 16 hours per simulation in a batch of 10 simulations per iteration of the outermost loop.
@@ 3.1 Discrete dislocation dynamics simulations
Evolution of a typical DDD simulation is illustrated in Fig. 2 at three stages: (a) instantiation of 1 m radius shear loops and prismatic loops as described in section 2.6, (b) evolution until a slip system reaches 0.2% strain difference from its initial state, and (c) final higher density configuration. These images correspond to the simulation measurements displayed in Fig. 3a.
@@
The prominence of rising densities is a consequence of our use of periodic boundaries. In a DDD convention in which plastic distortion of a shear loop resolved on its own slip system is always negative, applying a positive resolved shear stress causes loops within that slip system to shrink, contributing positively to the strain rate, while a negative will cause dislocation loops to grow and contribute negatively to the slip system’s strain rate. Thus the dislocation loops which contribute positively to strain rates are short-lived, shrinking until their inevitable self annihilation and often do not persist far beyond the 0.2% strain threshold. For this reason data from slip systems having primarily positive strain rates are not anticipated to be of imperative value to CP simulations, and the data sets produced by our simulations are acceptably sparse in the regions of positive strain rates. Fig. 3c demonstrates an occasional example counter to this trend, in which the applied load and configuration led to a slow shrinking of the majority of dislocation loops, specifically those on slip systems numbered 2 and 3, while the dislocation density on slip system 1 appears to have been stabilized under the load.
@@
Fig. 3 illustrates several other notable features found within our measurements. Deformation within the simulation presented in the second column is dominated by the single slip system numbered 22 whose dislocation density and plastic deformation show consistent, if not accelerating, growth. This is in contrast to the simulation of the first column in which the dislocation densities of slip systems 1 and 11, which have identical Burgers’ vectors but differing glide planes, are growing at rates which vary in time, presumably due to interactions between dislocations on the differing slip systems. It could be argued that the initially high strain rate of the first slip system was due to lack of interaction with the other slip systems resulting in a period of free movement which was soon to be suppressed when interactions became prominent, producing varying strain rates and noise in the functional relationship . The leading theory of how slip system interactions reduce strain rates is via the formation of junctions between dislocations whose mobilities are reduced or negligible [bulatov2006dislocation, PhysRevLett.121.085501]. In this particular example the two slip systems which appear to be interacting cannot form junctions because they have identical Burgers’ vectors. It should be noted that the specific configurational information which induces these interactions and the resulting strain rates is lost when the configuration is represented as dislocation densities per slip system, but could be recovered from the raw logs of the DDD simulations. The significant variations in strain rate measurements for a given stress tensor and list of densities are then represented by the fitted noise hyperparameter of the GP model.
The non-static nature of dislocation densities produced by the dynamics described above results in differences between the measured model domain values and those desired by the acquisition function. Generally, the longer a DDD runs, the farther from the targeted densities the system wanders. The resulting differences between targets and measurements are shown concretely by comparing Figures 6(b,e) or (e,f) in the Appendix. Additionally, the lack of precise control over the evolution of dislocation densities in DDD simulations prohibit deployment of a heteroscedastic noise model by impeding the required measurements of noise through resampling of identical domain values. Note however that the GP model is trained using the actual density and stress values measured during the DDD simulations, not to the target values. While the drift from the target values somewhat diminishes the accuracy of the BO at targeting specific regions of configuration space, it does not introduce a bias in the model.
@@ 3.2.1 Convergence on the model scale to test data
The non-negligible magnitude of the noise in DDD strain-rate measurements induces a lower limit on the error of our model, forcing the average error of its predictions to be at least as bad as the noise. This is shown by the similarity between the curves labeled ‘AVG of STDEV of test data’ and ‘AVG error of predictive mean’ in Figure 4a. This echos observations by another study[AKHONDZADEH2020104152] which discussed the effect of noise on fitted coefficients of the Kocks-Mecking model using a large database of DDD simulations. We concur that methods of reducing and compensating for the noise obtained in these measurements are needed to improve model predictions. While optimizing the reduced dimensional representation of all 24 slip system densities could possibly reduce noise, e.g. forest[PhysRevLett.89.255508], coplanar[MA20043603, NGUYEN2021102940], or glissile/sessile, our choice of a simpler specification avoided further complicating the acquisition function.
@@ 3.2.2 Evolution of the model’s physical meaningfulness
Due to the inverse relationship between feature sensitivity and hyperparameter values, fluctuations in the larger valued hyperparameters have lesser impact on model behavior due to their correspondence with features to which the model is least sensitive. Although the larger valued hyperparameters do not tightly converge to final values, their fluctuations have the least impact, while the most sensitive features have hyperparameters which are consistently so small that they are difficult to distinguish when presented on the same scale.
@@ Appendix A Distributions of simulation targets and eventual measurements
In Section 2.4 we described how our acquisition function was designed to address the differences in dimensionality between model domain and simulation parameters, while in Sections 2.5 and 2.7, we described simulation constraints and transformations, respectively, intended to improve the distributions of data presented to the model in the spirit of regularization so that it may produce a better fit. To assess the influence of these designs, transformations, and constraints we present here a study of the intrinsic distributions they collectively impart. We compare these intrinsic distributions to the final collection of training data obtained from measurements to gain insight into the ability of our BO implementation to navigate parameter space under the influence of the intrinsic distributions.
@@
The distribution of BO targets for particular slip system densities resemble neither distribution of the first or the remaining slip systems particularly well, indicating that the acquisition function had greater liberty in guiding slip system density specifications. However, the distribution of measured slip system density values which evolved during the simulations more closely resemble the distribution labeled “random, all but first slip system,” with the exception of the presence of a large number of slip systems having dislocation density. So while the acquisition function had the most liberty to determine dislocation density targets, their evolution during the simulation caused measurements to drift significantly far from their targets.